nikkixmarie:

A ball in the air has a height of h(t) + 16t^2 + 50t +3 feet after t seconds. a. what is the average rate of change for h(t) between t=0 and t=2? b. This means, on average, the height of the ball is (increasing or decreasing) for 0

Vocaloid:

rate of change formula [h(2) - h(0)] / (2-0) = ?

nikkixmarie:

sorry i dont know how to solve that

Vocaloid:

it's like our pizza problem

Vocaloid:

look at the formula h(0) = ?

nikkixmarie:

0?

Vocaloid:

h(t) + 16t^2 + 50t +3 h(0) = 16(0)^2 + 50(0) + 3 = ?

nikkixmarie:

0=3

Vocaloid:

h(0) = 3, good now find h(2)

nikkixmarie:

167?

Vocaloid:

good now find [h(2) - h(0)] / (2-0) using the values we obtained

nikkixmarie:

82?

Vocaloid:

yeah that's right, 82 m/s

Vocaloid:

so the height is increasing

nikkixmarie:

ok it has a drop down menu for part a: 82 (feet or seconds) per (foot or second)

nikkixmarie:

what do i choose

Vocaloid:

oh right it's in feet not meters sorry

Vocaloid:

feet per second

nikkixmarie:

thank you so much